OpenStudy (anonymous):

how would you add 6/25 and 1/2 without a calculator

OpenStudy (anonymous):

I cant find a common denominator

OpenStudy (anonymous):

\[\frac{a}{b}+\frac{c}{d}\text{=}\frac{a d+b c}{b d}\]

OpenStudy (anonymous):

so do I just plug in those numbers into the equation u just gave me?

OpenStudy (anonymous):

\[\frac{6*2+25*1}{25*2}=\frac{37}{50} \]

OpenStudy (anonymous):

a common denominator would be 50. so 12/50 + 25/50 = 37/50

OpenStudy (anonymous):

how do I find the domain of a graph, like what should I look for to find the two numbers for the domain, im confused